OpenStudy (anonymous):

One number is five more than another number and the sum of the two numbers is 55, what are the numbers?

OpenStudy (anonymous):

Let one number be x. So, other number is x+5 x+x+5=55 2x+5=55 2x=55-5 =50 x=25 x+5=30

OpenStudy (anonymous):

THNX

OpenStudy (anonymous):

Solve the following for x and y: {x = y + 5, x + y = 55 } x = 30, y = 25